# Multi-Knowledge-oriented Nighttime Haze Imaging Enhancer for Vision-driven Intelligent Transportation Systems

### v3

@@ Abstract
Salient object detection (SOD) plays a critical role in intelligent transportation systems (ITS), facilitating the detection and segmentation of key visual elements in an image. However, adverse imaging conditions such as haze during the day, low light, and haze at night severely degrade image quality and hinder reliable object detection in real-world scenarios. To address these challenges, we propose a multi-knowledge-oriented nighttime haze imaging enhancer (MKoIE), which integrates three tasks: daytime dehazing, low-light enhancement, and nighttime dehazing. The MKoIE incorporates two key innovative components: First, the network employs a task-oriented node learning mechanism to handle three specific degradation types: day-time haze, low light, and night-time haze conditions, with an embedded self-attention module enhancing its performance in nighttime imaging. In addition, multi-receptive field enhancement module that efficiently extracts multi-scale features through three parallel depthwise separable convolution branches with different dilation rates, capturing comprehensive spatial information with minimal computational overhead to meet the requirements of real-time ITS deployment. To ensure optimal image reconstruction quality and visual characteristics, we suggest a hybrid loss function. Extensive experiments on different types of weather/imaging conditions illustrate that MKoIE surpasses existing methods, enhancing the reliability, accuracy, and operational efficiency of ITS. The code is available at https://github.com/Ai-Chen-Lab/MKoIE.
Index Terms:
@@
Adverse weather, multi-knowledge learning, nighttime haze, image restoration, intelligent transportation systems (ITS)
I Introduction
@@
Salient object detection (SOD) [1, 2] is a critical technology in vision-driven intelligent transportation systems (VITS), enabling the rapid and accurate identification of key objects from real-time traffic images, such as vehicles, pedestrians, and road infrastructure. High-quality SOD is essential for ensuring safety, optimizing traffic flow, and supporting automated decision-making in modern transportation networks. However, as shown in Fig. 1, adverse imaging environmental conditions, such as low light, daytime or nighttime haze, can significantly affect the accuracy of SOD. Specifically, optical scattering from haze particles not only reduces contrast and sharpness but also distorts the spectral content of the captured scene [3]. Reduced illumination further exacerbates this issue by limiting signal strength, necessitating longer exposures or higher sensor gains, which result in additional electronic noise [4]. These environmental factors necessitate robust VITS that can effectively model and compensate for light attenuation, atmospheric scattering, and sensor-induced artifacts while maintaining measurement fidelity.
@@
Figure 1: The upper and lower sections illustrate daytime imaging scenarios, where varying haze densities impact object visibility. SOD operates on haze devices as a general-purpose preprocessing module, filtering salient data locally and transmitting it to the cloud for VITS.
@@
To address these challenges, researchers are studying non-learning and learning methods to improve the quality of degraded images. Traditional non-learning image restoration methods predominantly rely on histogram equalization [5] and the Retinex methods [6, 7, 8], which enhance image quality by adjusting brightness and contrast. However, these methods frequently lead to noise or color distortion, which undermine the reliability of ITS applications. In the context of image dehazing, classical methods such as dark channel prior (DCP) -based methods [9, 10] are effective under daytime conditions but encounter limitations when confronted with complex nighttime scenes. To further enhance image restoration in nighttime hazy conditions, researchers have optimized and refined Retinex and DCP models and proposed new physically-based imaging models [11, 12, 13] that are specifically tailored to handle the complexities of nighttime haze scenarios.
@@
Learning-based methods are widely used in image restoration tasks due to their powerful feature learning capabilities. For example, convolutional neural network (CNN)-based methods [14] significantly enhance the quality of low-light or hazy images in various tasks by learning different features from large-scale datasets. The attention mechanism can dynamically focus on important features and enhance the extraction and processing of key information, thus performing effectively in low-level vision tasks [15, 16, 17, 18]. Generative adversarial networks are used for unsupervised learning tasks [19, 20], pushing the boundaries of image processing in low-light scenarios. Diffusion models [21] have achieved success in low-light image enhancement by producing high-quality images through an iterative denoising process. To combine the advantages of traditional methods and learning methods, model-guided learning methods [22, 23, 24] have gradually been used to restore degraded images and have achieved satisfactory visual results. However, they are primarily designed for one task and cannot simultaneously handle three types of tasks: image dehazing (ID), low-light image enhancement (LLIE) or nighttime haze image enhancement (NHIE), and do not offer a unified solution for the variable imaging conditions encountered in ITS.
@@
  • Extensive experiments demonstrate that MKoIE significantly outperforms existing methods in image quality improvement and is a fast and lightweight model that achieves higher accuracy and reliability in VITS.
The remainder of this paper is organized as follows. Degradation models for nighttime haze are given in Section III. MKoIE is detailedly described in Section IV. Experimental results and discussions are provided in Section V. Conclusions are drawn in Section VI.
II Related Work
@@
II-A Daytime Image Dehazing
@@
II-B Low-light Image Enhancement
@@
II-C Nighttime Image Dehazing
@@
III Physical Imaging Models
@@
It needs to be noted that in practical applications of VITS, Eq. 1 needs to be adjusted according to specific circumstances. Due to the complexity of light source, particularly at night, directly applying above formula may not be sufficiently accurate. The challenges posed by varying lighting conditions, haze, and sensor noise require a more customized approach. To address these issues, our method focuses on three key image restoration tasks in VITS: ID, LLIE, and NHIE. To improve model accuracy under these conditions, we propose enhancements by integrating an illumination model and a noise model tailored for different scenes, which consider the unique interplay of haze, lighting, and sensor constraints.
@@
IV MKoIE: Multi-Knowledge-oriented Imaging Enhancer
IV-A Network Architecture
@@
IV-B Residual learning block
@@
IV-C Task-oriented Node Learning
@@
IV-C1 Task-Specific Module
@@
IV-C2 Self Attention-guided Node Learning Module
@@
IV-D Multi-Receptive Field Enhancement Module
@@
IV-E Self Attention Module
@@
IV-F Hybrid Loss Function
@@ IV-F2 Perceptual Loss
TABLE I: Comparison of ID results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
TABLE II: Comparison of LLIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
TABLE III: Comparison of NHIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@ IV-F3 Total Loss
V Experiments and Discussions
In this section, we conduct extensive experiments to highlight the outstanding ID, LLIE, and NHIE performance of MKoIE. We first provide an overview of the datasets and explain the implementation details. To illustrate the effectiveness of MKoIE, we then conduct both quantitative and qualitative analyses, comparing it with various state-of-the-art methods using synthetic and real-world low-visibility images. Additionally, we perform ablation studies to evaluate the contribution of each proposed module. Finally, we analyze and compare MKoIE in the context of the saliency detection task, showing its potential to enhance system performance in real-world transportation scenarios.
V-A Datasets and Implementation Details
V-A1 Training and Testing Datasets
The training and testing datasets include realistic single image dehazing (RESIDE)-OTS [44], which incorporates depth information, and the composite degradation dataset (CDD) [45]. Based on the imaging model in Sec. III, we extract atmospheric light values according to real conditions to synthesize more realistic degraded images [22]. By creating degraded images that closely resemble real-world conditions, we address the limitation of insufficient paired training datasets, thus enhancing the network’s restoration performance. During training, paired images are segmented into $256\times 256$ patches to achieve data augmentation and improve the network’s generalization capability. To evaluate the inference capabilities of MKoIE, we use a variety of synthetic and real-world low-visibility images.
V-A2 Evaluation Metrics
@@
V-A3 Competitive Methods
To evaluate the effectiveness of the MKoIE, we conducted quantitative and qualitative comparative analysis with competing methods, including ID methods DCP [9], FFANet [15], SDD [8], CEEF [41], ROP+ [10], DehazeFormer [16], and AoSRNet [22]; LLIE methods LIME [7], SDD [8] , CEEF [41], EFINet [42], ROP+ [10], SMNet [14], and AoSRNet [22]; and NHIE methods IENHC [11], GMLC [12], MRP [13], OFSD [43], 4KDehazing [20], CEEF [41] and ECNHI [23]. To ensure fairness and objectivity in the experiment, all methods are based on the source code published by the authors, and multiple experimental comparisons are carried out under the identical conditions to minimize the impact of extraneous factors on the results.
V-A4 Experiment Platform
@@
V-B Synthetic Degradation Analysis
V-B1 Dehazing
@@
Table I presents the comparative performance of MKoIE against seven ID methods on the RESIDE and CDD datasets. MKoIE achieved the highest average PSNR and SSIM scores, and the second best NIQE score, demonstrating its superiority in quantitative evaluation. Furthermore, MKoIE exhibits substantial advantages in qualitative comparisons. Fig. 5 illustrates the results of various methods on synthetic hazy images. The traditional DCP algorithm enhances contrast and removes some haze but often introduces color distortions and over-dehazes. Deep learning-based methods, such as FFANet and AoSRNet, improve haze removal and detail recovery but exhibit limitations when handling distant objects in natural scenes. DehazeFormer frequently produces prominent artifacts, particularly in sky regions, while CEEF, though capable of partial restoration, suffers from local distortions and loss of detail. Notably, MKoIE, despite being trained on diverse low-visibility images, demonstrates competitive performance comparable to state-of-the-art daytime ID methods, effectively addressing the challenges posed by synthetic and real-world hazy scenarios.
V-B2 Low-light Enhancement
@@
Table II and Fig. 5 illustrate that MKoIE significantly outperforms other methods in LLIE. Specifically, MKoIE demonstrates superior performance in image clarity and detail restoration, producing images with higher fidelity in terms of structural integrity while effectively reducing blurring and artifacts. LIME enhances visibility with noticeable improvements in color contrast and haze removal but introduces slight color distortions. SDD, while improving saturation in building areas and achieving a warmer appearance, suffers from blur and detail loss, particularly at the edges of buildings and within crowded regions. Both SMNet and EFINet exhibit over-sharpening effects, causing certain image details to appear excessively pronounced. In contrast, MKoIE generates images that closely resemble real-world scenes, with textures and colors more consistently preserved, thereby achieving a balance between enhancement and realism.
V-B3 Nighttime Haze Enhancement
@@
The evaluation of MKoIE was conducted using the RESIDE and CDD datasets, consistent with those in daytime ID and LLIE tasks. As shown in Fig. 5, while IENHC and ECNHI partially remove haze, residual haze remains visible, particularly in the background regions. In contrast, MKoIE cexhibits exceptional performance in color restoration, producing images with a natural tone devoid of oversaturation or color distortion, thereby enhancing both visual comfort and scene realism. In addition, MKoIE demonstrates superior detail preservation, with ground textures in the foreground and building outlines in the background clearly discernible. Quantitative evaluations in Table III further substantiates MKoIE’s superiority, as it achieves the best results in all metrics, underscoring its outstanding recovery performance under low-light and hazy conditions. Fig. 6’s histogram shows that our method achieves a balanced intensity distribution, along with superior brightness, color, and contrast enhancements. MKoIE produces a broad, well-spread histogram, indicating an optimal use of the intensity range. This corresponds to better brightness, more natural color distribution, and higher contrast compared to the other methods. Similarly, we conducted a detail enhancement comparison experiment, as shown in Fig. 7. The original image contains significant haze, especially around the object. IENHC yields a moderate clarity improvement: text becomes more legible, but background details remain largely unrecovered. MRP and OFSD both reduce haze and enhance image contrast and color saturation, though their results exhibit a slight remaining color cast. CEEF achieves only modest haze removal with a mild increase in contrast. In contrast, MKoIE shows pronounced improvements in overall color, contrast, and brightness. It also excels at restoring fine details, as evidenced by the clear textures visible in the red-boxed zoomed region in Figure 7. These visual enhancements correspond with MKoIE’s top quantitative metrics in Table III, confirming the method’s superior performance across both objective and subjective evaluations.
V-C Real-world Degradation Analysis

### v4

# Multi-Knowledge-oriented Nighttime Haze Imaging Enhancer for Vision-driven Intelligent Systems

@@ Abstract
Salient object detection (SOD) plays a critical role in Intelligent Imaging, facilitating the detection and segmentation of key visual elements in an image. However, adverse imaging conditions such as haze during the day, low light, and haze at night severely degrade image quality and hinder reliable object detection in real-world scenarios. To address these challenges, we propose a multi-knowledge-oriented nighttime haze imaging enhancer (MKoIE), which integrates three tasks: daytime dehazing, low-light enhancement, and nighttime dehazing. The MKoIE incorporates two key innovative components: First, the network employs a task-oriented node learning mechanism to handle three specific degradation types: day-time haze, low light, and night-time haze conditions, with an embedded self-attention module enhancing its performance in nighttime imaging. In addition, multi-receptive field enhancement module that efficiently extracts multi-scale features through three parallel depthwise separable convolution branches with different dilation rates, capturing comprehensive spatial information with minimal computational overhead to meet the requirements of real-time imaging deployment. To ensure optimal image reconstruction quality and visual characteristics, we suggest a hybrid loss function. Extensive experiments on different types of weather/imaging conditions illustrate that MKoIE surpasses existing methods, enhancing the reliability, accuracy, and operational efficiency of intelligent imaging.
keywords:
@@
Adverse weather, multi-knowledge learning, nighttime haze, image restoration, Intelligent imaging
journal: Applied Soft Computing
[…]
[1]organization=School of Computer Science and Engineering, addressline=University of Electronic Science and Technology of China, city=Chengdu, postcode=611731, state=Sichuan, country=China
[…]
[2]organization=Department of Logistics and Maritime Studies, addressline=The Hong Kong Polytechnic University, state=Hong Kong, country=China
[…]
[3]addressline=Chengdu Union Big Data Tech Inc., city=Chengdu, state=Sichuan, country=China
{highlights}
MKoIE is proposed to enhance imaging in low light, daytime haze, and nighttime haze.
Uses task-specific nodes and shared / separated networks for feature extraction.
SA and MRFE can enhance TNL’s ability to focus on critical regions and features.
A high-performance model, processes 2560×1440 images in 1.68s on a single GPU.
Extensive experiments show that the superiority of our MKoIE.
1 Introduction
@@
Salient object detection (SOD) [1] is a critical technology in vision-driven intelligent systems (VIS), enabling the rapid and accurate identification of key objects from real-time traffic images, such as vehicles, pedestrians, and buildings. High-quality SOD is essential for enhancing visual perception, enabling robust scene understanding, and supporting autonomous decision-making in VIS. However, as shown in Fig. 1, adverse imaging environmental conditions, such as low light, daytime or nighttime haze, can significantly affect the accuracy of SOD. Specifically, optical scattering from haze particles not only reduces contrast and sharpness but also distorts the spectral content of the captured scene [2]. Reduced illumination further exacerbates this issue by limiting signal strength, necessitating longer exposures or higher sensor gains, which result in additional electronic noise [3]. These environmental factors necessitate robust VIS that can effectively model and compensate for light attenuation, atmospheric scattering, and sensor-induced artifacts while maintaining measurement fidelity.
@@
Figure 1: The upper and lower sections illustrate daytime imaging scenarios, where varying haze densities impact object visibility. SOD operates on haze devices as a general-purpose preprocessing module, filtering salient data locally and transmitting it to the cloud for VIS.
@@
To address these challenges, researchers are studying non-learning and learning methods to improve the quality of degraded images. Traditional non-learning image restoration methods predominantly rely on histogram equalization [4] and the Retinex methods [5, 6, 7], which enhance image quality by adjusting brightness and contrast. However, these methods frequently lead to noise or color distortion, which undermine the reliability of VIS applications. In the context of image dehazing, classical methods such as dark channel prior (DCP) -based methods [8, 9] are effective under daytime conditions but encounter limitations when confronted with complex nighttime scenes. To further enhance image restoration in nighttime hazy conditions, researchers have optimized and refined Retinex and DCP models and proposed new physically-based imaging models [10, 11, 12] that are specifically tailored to handle the complexities of nighttime haze scenarios.
@@
Learning-based methods are widely used in image restoration tasks due to their powerful feature learning capabilities. For example, convolutional neural network (CNN)-based methods [13] significantly enhance the quality of low-light or hazy images in various tasks by learning different features from large-scale datasets. The attention mechanism can dynamically focus on important features and enhance the extraction and processing of key information, thus performing effectively in low-level vision tasks [14, 15, 16, 17]. Generative adversarial networks are used for unsupervised learning tasks [18, 19], pushing the boundaries of image processing in low-light scenarios. Diffusion models [20, 21] have achieved success in low-light image enhancement by producing high-quality images through an iterative denoising process. To combine the advantages of traditional methods and learning methods, model-guided learning methods [22, 23, 24] have gradually been used to restore degraded images and have achieved satisfactory visual results. However, they are primarily designed for one task and cannot simultaneously handle three types of tasks: image dehazing (ID), low-light image enhancement (LLIE) or nighttime haze image enhancement (NHIE), and do not offer a unified solution for the variable imaging conditions encountered in VIS.
@@
  • 3.
    Extensive experiments demonstrate that MKoIE significantly outperforms existing methods in image quality improvement and is a fast and lightweight model that achieves higher accuracy and reliability in VIS.
The remainder of this paper is organized as follows. Degradation models for nighttime haze are given in Section 3. MKoIE is detailedly described in Section 4. Experimental results and discussions are provided in Section 5. Conclusions are drawn in Section 6.
2 Related Work
@@
2.1 Daytime Image Dehazing
@@
2.2 Low-light Image Enhancement
@@
2.3 Nighttime Image Dehazing
@@
3 Physical Imaging Models
@@
It needs to be noted that in practical applications of VIS, Eq. 1 needs to be adjusted according to specific circumstances. Due to the complexity of light source, particularly at night, directly applying above formula may not be sufficiently accurate. The challenges posed by varying lighting conditions, haze, and sensor noise require a more customized approach. To address these issues, our method focuses on three key image restoration tasks in VIS: ID, LLIE, and NHIE. To improve model accuracy under these conditions, we propose enhancements by integrating an illumination model and a noise model tailored for different scenes, which consider the unique interplay of haze, lighting, and sensor constraints.
@@
4 MKoIE: Multi-Knowledge-oriented Imaging Enhancer
4.1 Network Architecture
@@
4.2 Residual learning block
@@
4.3 Task-oriented Node Learning
@@
4.3.1 Task-Specific Module
@@
4.3.2 Self Attention-guided Node Learning Module
@@
4.4 Multi-Receptive Field Enhancement Module
@@
4.5 Self Attention Module
@@
4.6 Hybrid Loss Function
@@ 4.6.2 Perceptual Loss
Table 1: Comparison of ID results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
Table 2: Comparison of LLIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@
Table 3: Comparison of NHIE results (mean $\pm$ std) with PSNR, SSIM, and NIQE on RESIDE and CDD datasets. The best results are in bold, and the second best are with underline.
@@ 4.6.3 Total Loss
5 Experiments and Discussions
In this section, we conduct extensive experiments to highlight the outstanding ID, LLIE, and NHIE performance of MKoIE. We first provide an overview of the datasets and explain the implementation details. To illustrate the effectiveness of MKoIE, we then conduct both quantitative and qualitative analyses, comparing it with various state-of-the-art methods using synthetic and real-world low-visibility images. Additionally, we perform ablation studies to evaluate the contribution of each proposed module. Finally, we analyze and compare MKoIE in the context of the saliency detection task, showing its potential to enhance system performance in real-world scenarios.
5.1 Datasets and Implementation Details
5.1.1 Training and Testing Datasets
The training and testing datasets include realistic single image dehazing (RESIDE)-OTS [48], which incorporates depth information, and the composite degradation dataset (CDD) [49]. Based on the imaging model in Sec. 3, we extract atmospheric light values according to real conditions to synthesize more realistic degraded images [22]. By creating degraded images that closely resemble real-world conditions, we address the limitation of insufficient paired training datasets, thus enhancing the network’s restoration performance. During training, paired images are segmented into $256\times 256$ patches to achieve data augmentation and improve the network’s generalization capability. To evaluate the inference capabilities of MKoIE, we use a variety of synthetic and real-world low-visibility images.
5.1.2 Evaluation Metrics
@@
5.1.3 Competitive Methods
To evaluate the effectiveness of the MKoIE, we conducted quantitative and qualitative comparative analysis with competing methods, including ID methods DCP [8], FFANet [14], SDD [7], CEEF [41], ROP+ [9], DehazeFormer [16], AoSRNet [22], GridFormer [42], and TaylorFormerV2 [43]; LLIE methods LIME [6], SDD [7] , CEEF [41], EFINet [44], ROP+ [9], SMNet [13], Reti-Diff[45], AnlightenDiff[46], and AoSRNet [22]; and NHIE methods IENHC [10], GMLC [11], MRP [12], OFSD [47], 4KDehazing [19], CEEF [41] and ECNHI [23]. To ensure fairness and objectivity in the experiment, all methods are based on the source code published by the authors, and multiple experimental comparisons are carried out under the identical conditions to minimize the impact of extraneous factors on the results.
5.1.4 Experiment Platform
@@
5.2 Synthetic Degradation Analysis
5.2.1 Dehazing
@@
Table 1 presents the comparative performance of MKoIE against nine ID methods on the RESIDE and CDD datasets. MKoIE achieved the highest average PSNR and SSIM scores, and the second best NIQE score, demonstrating its superiority in quantitative evaluation. Furthermore, MKoIE exhibits substantial advantages in qualitative comparisons. Fig. 5 illustrates the results of various methods on synthetic hazy images. The traditional DCP algorithm enhances contrast and removes some haze but often introduces color distortions and over-dehazes. Deep learning-based methods, such as FFANet and AoSRNet, improve haze removal and detail recovery but exhibit limitations when handling distant objects in natural scenes. TaylorFormer suffers from incomplete haze removal, particularly in background regions. GridFormer occasionally produces slight over-enhancement in contrast and saturation, which may lead to mild color shifts in sky regions. DehazeFormer frequently produces prominent artifacts, particularly in sky regions, while CEEF, though capable of partial restoration, suffers from local distortions and loss of detail. Notably, MKoIE, despite being trained on diverse low-visibility images, demonstrates competitive performance comparable to state-of-the-art daytime ID methods, effectively addressing the challenges posed by synthetic and real-world hazy scenarios.
5.2.2 Low-light Enhancement
@@
Table 2 and Fig. 5 illustrate that MKoIE significantly outperforms other methods in LLIE. Specifically, MKoIE demonstrates superior performance in image clarity and detail restoration, producing images with higher fidelity in terms of structural integrity while effectively reducing blurring and artifacts. LIME enhances visibility with noticeable improvements in color contrast and haze removal but introduces slight color distortions. SDD, while improving saturation in building areas and achieving a warmer appearance, suffers from blur and detail loss, particularly at the edges of buildings and within crowded regions. Both SMNet and EFINet exhibit over-sharpening effects, causing certain image details to appear excessively pronounced. Reti-Diff over-darkens the shadow areas, resulting in loss of fine detail.AnlightenDiff appears muted and under-saturated, reducing the perceptual quality of restoration. In contrast, MKoIE generates images that closely resemble real-world scenes, with textures and colors more consistently preserved, thereby achieving a balance between enhancement and realism.
5.2.3 Nighttime Haze Enhancement
@@
The evaluation of MKoIE was conducted using the RESIDE and CDD datasets, consistent with those in daytime ID and LLIE tasks. Since there are few NHIE methods, we only compared 7 methods. As shown in Fig. 5, while IENHC and ECNHI partially remove haze, residual haze remains visible, particularly in the background regions. In contrast, MKoIE cexhibits exceptional performance in color restoration, producing images with a natural tone devoid of oversaturation or color distortion, thereby enhancing both visual comfort and scene realism. In addition, MKoIE demonstrates superior detail preservation, with ground textures in the foreground and building outlines in the background clearly discernible. Quantitative evaluations in Table 3 further substantiates MKoIE’s superiority, as it achieves the best results in all metrics, underscoring its outstanding recovery performance under low-light and hazy conditions. Fig. 6’s histogram shows that our method achieves a balanced intensity distribution, along with superior brightness, color, and contrast enhancements. MKoIE produces a broad, well-spread histogram, indicating an optimal use of the intensity range. This corresponds to better brightness, more natural color distribution, and higher contrast compared to the other methods. Similarly, we conducted a detail enhancement comparison experiment, as shown in Fig. 7. The original image contains significant haze, especially around the object. IENHC yields a moderate clarity improvement: text becomes more legible, but background details remain largely unrecovered. MRP and OFSD both reduce haze and enhance image contrast and color saturation, though their results exhibit a slight remaining color cast. CEEF achieves only modest haze removal with a mild increase in contrast. In contrast, MKoIE shows pronounced improvements in overall color, contrast, and brightness. It also excels at restoring fine details, as evidenced by the clear textures visible in the red-boxed zoomed region in Figure 7. These visual enhancements correspond with MKoIE’s top quantitative metrics in Table 3, confirming the method’s superior performance across both objective and subjective evaluations.
5.3 Real-world Degradation Analysis